OpenStudy (bekkah323):

a/a^2-1 plus 2a/a^2-1 Equation below

OpenStudy (bekkah323):

\[\frac{ a }{ a^2-1 } + \frac{ 2a }{ a^2-a}\]

jimthompson5910 (jim_thompson5910):

the denominators are both equal to a^2-1 so you can combine the numerators and place that result over the common denominator

OpenStudy (bekkah323):

but one is a and one is 1?

jimthompson5910 (jim_thompson5910):

the numerators are _____ and _____ fill in the blanks

OpenStudy (bekkah323):

a and 2a

jimthompson5910 (jim_thompson5910):

add them

OpenStudy (bekkah323):

3a?

jimthompson5910 (jim_thompson5910):

we can do this because the denominators are the same if they weren't the same, then we cannot add them (yet)

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

so the final answer is \[\large \frac{3a}{a^2 - 1}\]

OpenStudy (bekkah323):

i don't get why the denominators are the same because one is a^2-1 and the other is a^2-a

jimthompson5910 (jim_thompson5910):

oh so that's a typo at the top then

jimthompson5910 (jim_thompson5910):

you have a^2 - 1 for both at the very top

jimthompson5910 (jim_thompson5910):

so the expression is \[\large \frac{a}{a^2 - 1} + \frac{2a}{a^2 - a}\] right?

OpenStudy (bekkah323):

yah i fixed it in the actual equation i wrote down, sorry?

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (bekkah323):

yes

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

the first step is to factor each denominator

jimthompson5910 (jim_thompson5910):

\[\large \frac{a}{a^2 - 1} + \frac{2a}{a^2 - a}\] \[\large \frac{a}{(a-1)(a+1)} + \frac{2a}{a(a-1)}\]

jimthompson5910 (jim_thompson5910):

the next step is to identify the LCD do you see what it is?

OpenStudy (bekkah323):

they both have an a-1 but the right side needs an a+1 and the left side needs an a?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

the LCD is a(a-1)(a+1)

jimthompson5910 (jim_thompson5910):

the first fraction needs an 'a' in the denominator, so you have to multiply top and bottom by 'a' to get \[\large \frac{a}{(a-1)(a+1)} + \frac{2a}{a(a-1)}\] \[\large \frac{a*a}{a(a-1)(a+1)} + \frac{2a}{a(a-1)}\] \[\large \frac{a^2}{a(a-1)(a+1)} + \frac{2a}{a(a-1)}\]

jimthompson5910 (jim_thompson5910):

and the second fraction needs (a+1) in the denominator like you said earlier, so multiply top and bottom of the second fraction by (a+1) to get.... \[\large \frac{a^2}{a(a-1)(a+1)} + \frac{2a}{a(a-1)}\] \[\large \frac{a^2}{a(a-1)(a+1)} + \frac{2a(a+1)}{a(a-1)(a+1)}\] \[\large \frac{a^2}{a(a-1)(a+1)} + \frac{2a^2+2a}{a(a-1)(a+1)}\]

OpenStudy (bekkah323):

ok so would the answer be\[\frac{ 3a^2 +2a}{ a(a-1)(a+1) }\]

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (bekkah323):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome